# Teaching Language Models to Evolve with Users: Dynamic Profile Modeling for Personalized Alignment

### v1

@@ Abstract
Personalized alignment is essential for enabling large language models (LLMs) to engage effectively in user-centric dialogue. While recent prompt-based and offline optimization methods offer preliminary solutions, they fall short in cold-start scenarios and long-term personalization due to their inherently static and shallow designs. In this work, we introduce the Reinforcement Learning for Personalized Alignment (RLPA) framework, in which an LLM interacts with a simulated user model to iteratively infer and refine user profiles through dialogue. The training process is guided by a dual-level reward structure: the Profile Reward encourages accurate construction of user representations, while the Response Reward incentivizes generation of responses consistent with the inferred profile. We instantiate RLPA by fine-tuning Qwen-2.5-3B-Instruct, resulting in Qwen-RLPA, which achieves state-of-the-art performance in personalized dialogue. Empirical evaluations demonstrate that Qwen-RLPA consistently outperforms prompting and offline fine-tuning baselines, and even surpasses advanced commercial models such as Claude-3.5 and GPT-4o. Further analysis highlights Qwen-RLPA’s robustness in reconciling conflicting user preferences, sustaining long-term personalization and delivering more efficient inference compared to recent reasoning-focused LLMs. These results emphasize the potential of dynamic profile inference as a more effective paradigm for building personalized dialogue systems.
@@ 1 Introduction
In recent years, aligning large language models (LLMs) with human values and intentions has become a crucial prerequisite for deploying them in interactive applications (Askell et al., 2021; Bai et al., 2022; Zhao et al., 2023). Alignment techniques—such as instruction tuning and reinforcement learning from human feedback (RLHF)—have significantly improved the helpfulness and harmlessness of model outputs by aligning them with broadly shared human preferences (Ouyang et al., 2022; Ji et al., 2023; Shen et al., 2023). This form of general alignment enables LLMs to follow instructions, avoid unsafe content, and exhibit socially acceptable behavior across a wide range of users.
However, such a one-size-fits-all alignment paradigm fails to account for the diversity of individual user needs, goals, and interaction styles (Sorensen et al., 2024; Kirk et al., 2024a; Jiang et al., 2024). As a result, it limits the model’s capacity to engage in personalized communication, which is essential for domains such as long-term dialogue, personalized education, and user-centered decision support (Zhang et al., 2024a; Tseng et al., 2024; Wu et al., 2024; Salemi et al., 2024a; Liu et al., 2025).
While personalized alignment holds great promise for enhancing user experience, achieving it in practice remains a significant challenge. A core difficulty stems from the inherently dynamic and evolving nature of personalization (Shi et al., 2024; Jiang et al., 2025). In particular, effective personalized alignment requires the model to continuously infer and adapt to user-specific attributes—such as preferences, goals, and beliefs—throughout the interaction. In real-world settings, the model often encounters a cold-start scenario, where little or no prior user information is available. As a result, it must dynamically construct and refine user representations based solely on the dialogue context, enabling long-term adaptation and the development of coherent user profiles across extended conversations (Zhao et al., 2025a; Xie et al., 2025).
@@
While recent efforts have sought to address these challenges of personalization, they remain inadequate in meeting the demands of dynamic and adaptive user needs. As shown in Figure 1(a), prompt-based methods—such as profile-augmented prompting (Wang et al., 2023; Richardson et al., 2023; Pandey et al., 2024; Li et al., 2024a) and retrieval-augmented prompting (Zhang et al., 2024b; Li et al., 2024b; Salemi et al., 2024b; Zhuang et al., 2024)—typically provide only superficial personalization, relying on static templates to inject user-specific information. This approach limits both the flexibility and expressiveness of model outputs and hampers the integration of long-term user knowledge due to the constraints of the context window (Liu et al., 2025). Meanwhile, offline optimization methods in Figure 1(b), such as supervised fine-tuning (SFT) (Ouyang et al., 2022; Clarke et al., 2024; Tan et al., 2024; Peng et al., 2024) and direct preference optimization (DPO) (Rafailov et al., 2023; Jang et al., 2023; Kirk et al., 2024b; Zollo et al., 2024; Chen et al., 2025), require large-scale labeled datasets, making them impractical in cold-start scenarios where user data is scarce or unavailable. Moreover, these methods tend to generalize poorly across users due to their static nature (Xu et al., 2024; Lin et al., 2024; Chu et al., 2025), rendering them inflexible for real-time adaptation during interactions.
@@
We fine-tune the Qwen-2.5-3B-Instruct model (Yang et al., 2024) using our proposed RLPA framework, resulting in the Qwen-RLPA model. Experimental results show that Qwen-RLPA substantially outperforms both prompt-based and offline optimization baselines in terms of personalization quality. Notably, it surpasses leading proprietary systems such as Claude-3.5 and DeepSeek-V3, and achieves performance on par with GPT-4o. Further analysis indicates that Qwen-RLPA is capable of sustaining coherent, personalized responses throughout extended long-term dialogues, effectively resolving preference conflicts and dynamically adapting its behavior. Moreover, when compared against recent state-of-the-art reasoning LLMs, including DeepSeek-R1 (Guo et al., 2025) and OpenAI-o3 (OpenAI, 2025), Qwen-RLPA delivers superior performance with significantly greater inference efficiency. These results highlight the potential of dynamically constructed user profiles as a more appropriate and effective reasoning paradigm for personalized dialogue systems.
@@ 2.2 Reinforcement Learning for Policy Optimization
$J(\pi)=\mathbb{E}_{\pi}\left[\sum{t=1}^{T}\gamma^{t-1}R_{t}\right]$ (1)
@@ 3.1 Simulated User Design
Each simulated user is initialized with a given profile $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, where each $p_{i}$ denotes a user attribute such as preferences, personality traits, or goals. These attributes are embedded into the system prompt of the user model, which conditions its behavior throughout the dialogue (Brown et al., 2020; Kojima et al., 2022; Zhao et al., 2025b).
@@
To assess whether the user model satisfies the desired properties, we conduct human evaluations using various base models as user simulators. The results reveal notable differences in profile fidelity and behavioral consistency across models. Balancing performance with computational cost, we ultimately select GPT-4o-mini as the user model for all subsequent experiments. Comprehensive evaluation results and comparisons are provided in Appendix H.
@@ 3.2 Profile Reward Function Design
To enable structured inference and efficient matching, we represent the user profile $\mathcal{P}$ using a slot-value format, where each slot $p_{i}$ denotes a predefined attribute category paired with a specific value (see Figure 7 in Appendix D.2 for detailed examples). At each dialogue turn $t$, the model is trained to produce its current estimate $\hat{\mathcal{P}}_{t}={\hat{p}_{1,t},\dots,\hat{p}_{n,t}}$ based on the dialogue history up to that point. This structured representation supports explicit supervision by allowing direct comparison between the predicted and ground-truth profile slots. The reward is calculated using a slot-wise matching score:
@@ 3.3 Response Reward Function Design
$\small R=N\cdot R\cdot L\cdot G\cdot F$ (4)
@@
To select a reliable reward model, we benchmark several LLMs (GPT-4o, DeepSeek-V3) by comparing their scoring consistency with human judgments. GPT-4o yields the highest agreement, but for efficiency, we adopt GPT-4o-mini during RL training. Full results are shown in Appendix I.2.
@@ 3.4 Training with Proximal Policy Optimization (PPO)
To perform optimization in RLPA, we adopt Proximal Policy Optimization (PPO) (Schulman et al., 2017), a widely used policy gradient algorithm. At each dialogue turn $t$, the model receives a combined reward signal that supervises both user modeling and personalized response generation:
@@
This turn-level reward encourages the model to continuously infer, refine, and utilize user profiles throughout the interaction. The full PPO optimization objective are provided in Appendix C.
@@ 4.1 Experimental Setup
We instantiate our RLPA framework using the Qwen-2.5-3B-Instruct (Yang et al., 2024). For the user simulator, we adopt GPT-4o-mini, selected based on our human evaluation study (see Appendix H). The reward model is also implemented using GPT-4o-mini and prompted to assess response alignment with user profiles across four personalization dimensions. To facilitate profile supervision, we preprocess the ALOE training set by converting each user profile into a structured slot-value format, enabling fine-grained attribute tracking aligned with our reward design. The full construction procedure is detailed in Appendix D.1.
@@
We evaluate on the ALOE benchmark (Wu et al., 2025), which provides multi-turn dialogues annotated with user profiles across diverse attributes for personalized dialogue evaluation.
@@
We adopt the average alignment score (AVG.), normalized improvement ratio (N-IR) and normalized coefficient of determination (N-$R^{2}$). Details on metrics calculation and evaluation prompts are in Appendix D.3 and Appendix D.4, respectively.
@@
We compare RLPA against two major categories of baseline approaches for personalized alignment. Prompt-based Methods: (1) Reminder (Zhao et al., 2025a), (2) Self-Critic (Zhao et al., 2025a), (3) Chain-of-Thought (CoT) (Wei et al., 2022) and (4) RAG (Zhao et al., 2025a). Offline Optimization Methods: (5) Supervised Finetuning (SFT) (Ouyang et al., 2022) and (6) Direct Preference Optimization (DPO) (Rafailov et al., 2023). Please refer to Appendix E for the detailed description of the baseline methods.
@@
We implement our RLPA training pipeline using the OpenRLHF (Hu et al., 2024) and vLLM (Kwon et al., 2023) frameworks for scalable and stable reinforcement learning with LLMs. All experiments are conducted on 8 NVIDIA A100 80GB GPUs. For detailed hyper-parameter settings, please refer to Appendix F.
@@
[…]
@@ 5.1 Ablation Study
To assess the individual contributions of the Profile Reward (PR) and Response Reward (RR), we conduct ablation experiments by disabling one reward component at a time during training. Table 2 reports the alignment results across dialogue turns. The reward progression curves for each training setup are provided in Appendix J, indicating that training remains stable across all configurations, with smooth convergence patterns and no signs of reward collapse or instability.
@@ 5.4 Comparison with Reasoning LLMs
We consider inferring user profiles as a domain-specific form of reasoning in personalized dialogue. Accordingly, we compare our Qwen-RLPA model against several reasoning-centric LLMs, including DeepSeek-R1 (Guo et al., 2025), GPT-o3-mini (OpenAI, 2025), and QwQ-32B (Qwen, 2025).
@@ 6 Related Works
On the one hand, prompt-based methods, including profile-augmented prompting (Wang et al., 2023; Richardson et al., 2023; Pandey et al., 2024; Li et al., 2024a) and retrieval-augmented prompting (Zhang et al., 2024b; Li et al., 2024b; Salemi et al., 2024b; Zhuang et al., 2024), typically rely on static templates to inject user-specific information. While simple to implement, these methods offer only superficial personalization, are constrained by context length limitations (Liu et al., 2025), and lack mechanisms for long-term memory or adaptive behavior.
On the other hand, offline optimization approaches, such as supervised fine-tuning (SFT) (Ouyang et al., 2022; Clarke et al., 2024; Tan et al., 2024; Peng et al., 2024) and direct preference optimization (DPO) (Rafailov et al., 2023; Jang et al., 2023; Kirk et al., 2024b; Zollo et al., 2024; Chen et al., 2025), aim to train models to produce profile-consistent responses. However, they require extensive labeled data, making them unsuitable for cold-start scenarios. Moreover, their static nature limits generalization across diverse users and hinders real-time adaptation (Xu et al., 2024; Chu et al., 2025).
@@ 7 Conclusion
Appendix A Limitation and Future Works

### v2

@@ Abstract
Personalized alignment is essential for enabling large language models (LLMs) to engage effectively in user-centric dialogue. While recent prompt-based and offline optimization methods offer preliminary solutions, they fall short in cold-start scenarios and long-term personalization due to their inherently static and shallow designs. In this work, we introduce the Reinforcement Learning for Personalized Alignment (RLPA) framework, in which an LLM interacts with a simulated user model to iteratively infer and refine user profiles through dialogue. The training process is guided by a dual-level reward structure: the Profile Reward encourages accurate construction of user representations, while the Response Reward incentivizes generation of responses consistent with the inferred profile. We instantiate RLPA by fine-tuning Qwen-2.5-3B-Instruct, resulting in Qwen-RLPA, which achieves state-of-the-art performance in personalized dialogue. Empirical evaluations demonstrate that Qwen-RLPA consistently outperforms prompting and offline fine-tuning baselines, and even surpasses advanced commercial models such as Claude-3.5 and GPT-4o. Further analysis highlights Qwen-RLPA’s robustness in reconciling conflicting user preferences, sustaining long-term personalization and delivering more efficient inference compared to recent reasoning-focused LLMs. These results emphasize the potential of dynamic profile inference as a more effective paradigm for building personalized dialogue systems. Our code is available at: https://github.com/XingYuSSS/RLPA.
@@ 1 Introduction
In recent years, aligning large language models (LLMs) with human values and intentions has become a crucial prerequisite for deploying them in interactive applications (askell2021general; bai2022constitutional; zhao2023survey). Alignment techniques—such as instruction tuning and reinforcement learning from human feedback (RLHF)—have significantly improved the helpfulness and harmlessness of model outputs by aligning them with broadly shared human preferences (ouyang2022training; ji2023ai; shen2023large). This form of general alignment enables LLMs to follow instructions, avoid unsafe content, and exhibit socially acceptable behavior across a wide range of users.
However, such a one-size-fits-all alignment paradigm fails to account for the diversity of individual user needs, goals, and interaction styles (sorensen2024position; kirk2024benefits; jiang2024peek). As a result, it limits the model’s capacity to engage in personalized communication, which is essential for domains such as long-term dialogue, personalized education, and user-centered decision support (zhang2024personalization; tseng2024two; wu2024personalized; salemi2024lamp; liu2025survey).
While personalized alignment holds great promise for enhancing user experience, achieving it in practice remains a significant challenge. A core difficulty stems from the inherently dynamic and evolving nature of personalization (shi2024wildfeedback; jiang2025know). In particular, effective personalized alignment requires the model to continuously infer and adapt to user-specific attributes—such as preferences, goals, and beliefs—throughout the interaction. In real-world settings, the model often encounters a cold-start scenario, where little or no prior user information is available. As a result, it must dynamically construct and refine user representations based solely on the dialogue context, enabling long-term adaptation and the development of coherent user profiles across extended conversations (zhao2025llms; xie2025survey).
@@
While recent efforts have sought to address these challenges of personalization, they remain inadequate in meeting the demands of dynamic and adaptive user needs. As shown in Figure 1(a), prompt-based methods—such as profile-augmented prompting (wang2023chain; richardson2023integrating; pandey2024cos; li2024learning) and retrieval-augmented prompting (zhang2024llm; li2024hello; salemi2024optimization; zhuang2024hydra)—typically provide only superficial personalization, relying on static templates to inject user-specific information. This approach limits both the flexibility and expressiveness of model outputs and hampers the integration of long-term user knowledge due to the constraints of the context window (liu2025survey). Meanwhile, offline optimization methods in Figure 1(b), such as supervised fine-tuning (SFT) (ouyang2022training; clarke2024peft; tan2024personalized; peng2024pocketllm) and direct preference optimization (DPO) (rafailov2023direct; jang2023personalized; kirk2024prism; zollo2024personalllm; chen2025deeper), require large-scale labeled datasets, making them impractical in cold-start scenarios where user data is scarce or unavailable. Moreover, these methods tend to generalize poorly across users due to their static nature (xu2024dpo; lin2024limited; chu2025sft), rendering them inflexible for real-time adaptation during interactions.
@@
We fine-tune the Qwen-2.5-3B-Instruct model (yang2024qwen2) using our proposed RLPA framework, resulting in the Qwen-RLPA model. Experimental results show that Qwen-RLPA substantially outperforms both prompt-based and offline optimization baselines in terms of personalization quality. Notably, it surpasses leading proprietary systems such as Claude-3.5 and DeepSeek-V3, and achieves performance on par with GPT-4o. Further analysis indicates that Qwen-RLPA is capable of sustaining coherent, personalized responses throughout extended long-term dialogues, effectively resolving preference conflicts and dynamically adapting its behavior. Moreover, when compared against recent state-of-the-art reasoning LLMs, including DeepSeek-R1 (guo2025deepseek) and OpenAI-o3 (openai2025o3blog), Qwen-RLPA delivers superior performance with significantly greater inference efficiency. These results highlight the potential of dynamically constructed user profiles as a more appropriate and effective reasoning paradigm for personalized dialogue systems.
@@ 2.2 Reinforcement Learning for Policy Optimization
$J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}\gamma^{t-1}R_{t}\right]$ (1)
@@ 3.1 Simulated User Design
Each simulated user is initialized with a given profile $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, where each $p_{i}$ denotes a user attribute such as preferences, personality traits, or goals. These attributes are embedded into the system prompt of the user model, which conditions its behavior throughout the dialogue (brown2020language; kojima2022large; DBLP:conf/nips/ZhengWZN0C24; zhao2025beware).
@@
To assess whether the user model satisfies the desired properties, we conduct human evaluations using various base models as user simulators. The results reveal notable differences in profile fidelity and behavioral consistency across models. Balancing performance with computational cost, we ultimately select GPT-4o-mini as the user model for all subsequent experiments. Comprehensive evaluation results and comparisons are provided in Appendix G.
@@ 3.2 Profile Reward Function Design
To enable structured inference and efficient matching, we represent the user profile $\mathcal{P}$ using a slot-value format, where each slot $p_{i}$ denotes a predefined attribute category paired with a specific value (see Figure 7 in Appendix C.2 for detailed examples). At each dialogue turn $t$, the model is trained to produce its current estimate $\hat{\mathcal{P}}_{t}={\hat{p}_{1,t},\dots,\hat{p}_{n,t}}$ based on the dialogue history up to that point. This structured representation supports explicit supervision by allowing direct comparison between the predicted and ground-truth profile slots. The reward is calculated using a slot-wise matching score:
@@ 3.3 Response Reward Function Design
$\small R^{\text{response}}_{t}=N\cdot R\cdot L\cdot G\cdot F$ (4)
@@
To select a reliable reward model, we benchmark several LLMs (GPT-4o, DeepSeek-V3) by comparing their scoring consistency with human judgments. GPT-4o yields the highest agreement, but for efficiency, we adopt GPT-4o-mini during RL training. Full results are shown in Appendix H.2.
@@ 3.4 Training with Proximal Policy Optimization (PPO)
To perform optimization in RLPA, we adopt Proximal Policy Optimization (PPO) (schulman2017proximal), a widely used policy gradient algorithm. At each dialogue turn $t$, the model receives a combined reward signal that supervises both user modeling and personalized response generation:
@@
This turn-level reward encourages the model to continuously infer, refine, and utilize user profiles throughout the interaction. The full PPO optimization objective are provided in Appendix B.
@@ 4.1 Experimental Setup
We instantiate our RLPA framework using the Qwen-2.5-3B-Instruct (yang2024qwen2). For the user simulator, we adopt GPT-4o-mini, selected based on our human evaluation study (see Appendix G). The reward model is also implemented using GPT-4o-mini and prompted to assess response alignment with user profiles across four personalization dimensions. To facilitate profile supervision, we preprocess the ALOE training set by converting each user profile into a structured slot-value format, enabling fine-grained attribute tracking aligned with our reward design. The full construction procedure is detailed in Appendix C.1.
@@
We evaluate on the ALOE benchmark (wu2025aligning), which provides multi-turn dialogues annotated with user profiles across diverse attributes for personalized dialogue evaluation.
@@
We adopt the average alignment score (AVG.), normalized improvement ratio (N-IR) and normalized coefficient of determination (N-$R^{2}$). Details on metrics calculation and evaluation prompts are in Appendix C.3 and Appendix C.4, respectively.
@@
We compare RLPA against two major categories of baseline approaches for personalized alignment. Prompt-based Methods: (1) Reminder (zhao2025llms), (2) Self-Critic (zhao2025llms), (3) Chain-of-Thought (CoT) (wei2022chain) and (4) RAG (zhao2025llms). Offline Optimization Methods: (5) Supervised Finetuning (SFT) (ouyang2022training) and (6) Direct Preference Optimization (DPO) (rafailov2023direct). Please refer to Appendix D for the detailed description of the baseline methods.
@@
We implement our RLPA training pipeline using the OpenRLHF (hu2024openrlhf) and vLLM (kwon2023efficient) frameworks for scalable and stable reinforcement learning with LLMs. All experiments are conducted on 8 NVIDIA A100 80GB GPUs. For detailed hyper-parameter settings, please refer to Appendix E.
@@
Align. Score (AVG.) $\uparrow$ N-IR $\uparrow$ N-$R^{2}$ $\uparrow$ Align. Score (AVG.) $\uparrow$ N-IR $\uparrow$ N-$R^{2}$ $\uparrow$
@@ 5.1 Ablation Study
To assess the individual contributions of the Profile Reward (PR) and Response Reward (RR), we conduct ablation experiments by disabling one reward component at a time during training. Table 2 reports the alignment results across dialogue turns. The reward progression curves for each training setup are provided in Appendix I, indicating that training remains stable across all configurations, with smooth convergence patterns and no signs of reward collapse or instability.
@@ 5.4 Comparison with Reasoning LLMs
We consider inferring user profiles as a domain-specific form of reasoning in personalized dialogue. Accordingly, we compare our Qwen-RLPA model against several reasoning-centric LLMs, including DeepSeek-R1 (guo2025deepseek), GPT-o3-mini (openai2025o3blog), and QwQ-32B (qwen2025qwqblog).
@@ 6 Related Works
On the one hand, prompt-based methods, including profile-augmented prompting (wang2023chain; richardson2023integrating; pandey2024cos; li2024learning) and retrieval-augmented prompting (zhang2024llm; li2024hello; salemi2024optimization; zhuang2024hydra; qiu2025measuring), typically rely on static templates to inject user-specific information. While simple to implement, these methods offer only superficial personalization, are constrained by context length limitations (liu2025survey), and lack mechanisms for long-term memory or adaptive behavior.
On the other hand, offline optimization approaches, such as supervised fine-tuning (SFT) (ouyang2022training; clarke2024peft; tan2024personalized; peng2024pocketllm) and direct preference optimization (DPO) (rafailov2023direct; jang2023personalized; kirk2024prism; zollo2024personalllm; chen2025deeper), aim to train models to produce profile-consistent responses. However, they require extensive labeled data, making them unsuitable for cold-start scenarios. Moreover, their static nature limits generalization across diverse users and hinders real-time adaptation (xu2024dpo; chu2025sft).
@@ 7 Conclusion
8 Limitation and Future Works